OpenStudy (anonymous):

PLEASE HELP! FAN AND MEDAL What are the missing values in the following geometric sequence? __, 10, 5, __, __.

OpenStudy (anonymous):

Is this all the information given?

OpenStudy (anonymous):

yes and then the answers. I know I need to find the common ratio but i dont know how exactly

OpenStudy (anonymous):

What are the choices

OpenStudy (anonymous):

20, start fraction five over two end fraction comma start fraction five over four end fraction 15, 0, −5 20, 1, start fraction one over two end fraction −20, start fraction five over two end fraction comma negative start fraction five over four end fraction

OpenStudy (anonymous):

Okay so it is given: 10, 5 What could have done? 10/2=5?

OpenStudy (anonymous):

If so, then the first digit must have been 20

OpenStudy (anonymous):

okay but I thought I would of had to multiply

OpenStudy (anonymous):

so 20, 10, 5, 5/2, 5/4

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (anonymous):

You would have if the geometric sequence was going up.. but since 10 came before 5.. you divide by 2

OpenStudy (anonymous):

Oh that makes sense! thank you so much

OpenStudy (anonymous):

Two geometric means are inserted between 6 and 384 so the four numbers form a geometric sequence. What are these two numbers? Hint: an = a1(r)n − 1, where a1 is the first term and r is the common ratio.

OpenStudy (anonymous):

an=a1(r)^n-1

OpenStudy (anonymous):

Okay, sorry I was helping someone..

OpenStudy (anonymous):

So we have 6,___, ___, 384 right

OpenStudy (anonymous):

its okay! no rush :) I appreciate you helping me

OpenStudy (anonymous):

−24, 96 −24, −96 24, 96 22, 108 these are the answers

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i would divide?

OpenStudy (anonymous):

Alright.. simple.. this time we multiply because 6 is the first term and we have to get to 384.

OpenStudy (anonymous):

actually I take that back I would multiply because its going up

OpenStudy (anonymous):

you would multiply by 2?

OpenStudy (anonymous):

So from 6 what do we multiply to get to the second term and the same to the third term and to the last term? Try numbers from 1-9 and see...

OpenStudy (anonymous):

So 6 x (a number) = ans ans x (a number) = ans2 ans2 x (a number) = ans3 ans3 x (a number) = 384

OpenStudy (anonymous):

the (a number) must be the same

OpenStudy (anonymous):

Definitely not a 1, 2, 3 because you will get a last term lower than 384, right?

OpenStudy (anonymous):

Not 9, 8, 7, 6 because the last term will be too high.. try 4 and 5

OpenStudy (anonymous):

6(x), N(x), N(x), 384 6(4), 24(4), 96(4), 384 Match? 24 and 96